OpenStudy (anonymous):

An original rectangle has a length of 10 and a width of 28. What happens to the area if the new width is one quarter the original width? Verify your answer.

OpenStudy (anonymous):

The new width is going to be \[\frac{1}{4}(28) = 7\] So the area is now A = (10)(7) = 70, which happens to be 1/4 the old area

OpenStudy (anonymous):

\[10\times 28\] vs \[10\times 7\]

OpenStudy (anonymous):

first one is 280, second one is 70, one fourth of the first one what joemath said

OpenStudy (anonymous):

Specifically, in your case it sounds like your teacher wants you to intuit the fact that the area is reduced by a factor of four, then verify it with the above multiplication.